OpenStudy (anonymous):

Given the equation y − 4 = three fourths(x + 8) in point-slope form, identify the equation of the same line in slope-intercept form. −three fourthsx + y = 10 3x − 4y = −40 y = three fourthsx + 12 y = three fourthsx + 10

OpenStudy (anonymous):

@ash2326 ? :)

OpenStudy (anonymous):

@amistre64 ?

OpenStudy (anonymous):

@ParthKohli ?

OpenStudy (amistre64):

in other words, solve for y \[y-a=m(x-b)\] \[y-a=mx-mb\] \[y=mx-mb+a\]

OpenStudy (anonymous):

y - 4 = 3/4(x + 8) -- distribute the 3/4 through the parenthesis) y - 4 = 3/4x + 6 (you want to get y by itself, so add 4 to both sides) y - 4 + 4 = 3/4x + 6 + 4 (combine like terms) y = 3/4x + 10 (slope intercept form) If you have any questions about what I have done, please ask. I want you to understand this .